OpenStudy (anonymous):

A plumbing contracter charges $40 per hour plus a call-out fee of $25. if the contracter charged $165 for a job, how long did the job take?

ganeshie8 (ganeshie8):

say it took him \(t\) hours

ganeshie8 (ganeshie8):

that means he must have charged \(40\times t\) amount

ganeshie8 (ganeshie8):

and an additional $25 makes it : \[40\times t + 25\] yes ?

ganeshie8 (ganeshie8):

set that equal to 165 and solve \(t\)

ganeshie8 (ganeshie8):

\[40\times t + 25 = 165\]

OpenStudy (anonymous):

So is the answer to this 165

OpenStudy (danjs):

Total Amount Charged = [Hourly Rate] x [#Hours] + [Call-out Fee]

OpenStudy (anonymous):

Do you know the answer?

OpenStudy (danjs):

I didnt calculate it, you need to put the numbers.

OpenStudy (anonymous):

idk where they go xD sorry im really dumb at maths

OpenStudy (danjs):

Total amount charged = 165 and call-out fee = 25 and hourly rate = 40

OpenStudy (danjs):

@ganeshie8 wrote the equation out for you above

OpenStudy (anonymous):

whats the t

OpenStudy (danjs):

so you have, 40*t + 25 = 165 What would you do first to solve for the amount of time t ?